As. Find a power-series expansion the General solution to the differentar equation yll-Sxy tyo
Answer #1

Let the power seves expansion g(n) = [annn n=0 be the general solution of the given differential equation. Now, differenti itNow the transformed differential eqn n-2 Ź n(n-1) anni - 5 n E hannn na2 + annn =0 n=0 E n(n-1) an yn-2 8W - 01 nannn ↑ na 2+60+2) (0+1) aq n° + [n+2)(n+1) 90+2 yn n=1 -52 nannn nanna taon+ ܐܶ܀ anna=o n=1 n=1 → E [{n+2) (n+1) ant2 - 5 nant an Inn n=i na = (n+2)(n+1) antz - (55-1) an = 0 9n+2 = 5n-1) ann21 (n+2)(n+1) Put n=1 5.11 1 44 (1+2)(1+1) 4 3.2 34 n=2, an 16-1 = 92If n is even. then (0-2) is also even then 5h-1 anta (1 +2)(n+1) 5 (1–2)-1.90-2 (n+2)(n+i) (n=2+2)(1+2+1) 5h-1 5 since on = 5

(n+2)(n+i) ((1-2)+2)(0-24) (2+2)(2+1) ant2 = le 51-1 5(n-2)-1 5.2-1 . ag If h is odd. then 1-2 is odot: then an+2= 5n-1 5(n-2Now, aanti nalt annel 92n+1+2-2 (5 (2n-1) -1) (6 (2003)-1) ( 20–4 +2) (2n-1 +1) (20-3+2) (20-3+1) le 92n-1) +2= 51-1 (1+2)(1+Now aan (21-2+ (ต-1)+) 1 .99 5 โ2(-]- ((n-1) +2) (2(ต-) +1) 52 - 1 (1+1) (241) Thus th *N aan + 12h+) () = 9thy h= 0 TT 5 (12

for a given a​​​​​​0  and a​​​​​​1 .